# Holistic Optimization Framework for FPGA Accelerators

### v2

@@ Abstract
Customized accelerators have transformed modern computing by delivering substantial gains in energy efficiency and performance through hardware specialization. Field-Programmable Gate Arrays play a critical role in this paradigm, offering unmatched flexibility and high-performance potential. High-Level Synthesis and source-to-source compilers have simplified Field-Programmable Gate Arrays design by translating high-level programming languages into hardware descriptions enriched with directives. However, achieving high Quality of Results remains a significant challenge, requiring intricate code transformations, strategic directive placement, and optimized data communication.
This paper presents Prometheus, a comprehensive framework that integrates key optimizations, including task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution, into a unified design space. By leveraging Non-Linear Problem methodologies, Prometheus explores this design space to deliver efficient solutions under strict resource constraints, facilitating automatic bitstream generation. The framework addresses limitations in existing methods and demonstrates superior performance, achieving high Quality of Results across complex computation-bound applications.
@@ I Introduction
The rise of customized accelerators has transformed modern computing, delivering substantial improvements in energy efficiency and performance through hardware specialization. Field-Programmable Gate Arrays (FPGAs) have emerged as a particularly versatile solution, valued for their flexibility and ability to achieve high performance across diverse applications. High-level synthesis (HLS) tools and source-to-source compilers [20, 80, 75, 32, 61, 49, 76, 31, 78, 37, 79, 42, 59, 9] have simplified hardware design by translating high-level programming languages (e.g., C++, Python) enriched with hardware-specific directives into synthesizable hardware descriptions. However, the effectiveness of these tools depends heavily on the quality of the input code and the strategic application of hardware directives, making it challenging to achieve high Quality of Results (QoR) without significant manual effort.
Despite advancements in the field, designing optimized FPGA solutions remains a complex task. Existing approaches [57, 58, 66, 65, 63, 25, 81, 65, 37, 78, 77, 59, 70, 9, 31] often address specific aspects of FPGA design in isolation, leaving gaps in areas such as efficient code transformation, effective concurrency management, and resource optimization. Moreover, challenges such as overlapping computation and communication, balancing intra-task parallelism, and managing FPGA resource constraints—including multi-Super Logic Region (SLR) architectures—further complicate the development process. These challenges are exacerbated by the lack of holistic frameworks that integrate multiple optimization strategies while ensuring synthesizable output.
Two commonly used paradigms—shared buffer and dataflow—have shown promise in improving FPGA performance. Shared buffers enable data reuse and promote parallelism by caching information on-chip, which is particularly beneficial for computation-bound kernels. However, their efficacy is often constrained by the limited on-chip memory, preventing full data caching and hindering the ability to overlap computation and communication effectively. On the other hand, the dataflow paradigm facilitates efficient data movement and task concurrency by leveraging FIFO (First-In-First-Out) communication. While this approach is effective for memory-bound kernels, it struggles to exploit shared memory and parallelism within tasks, and its independent task compilation can limit resource sharing opportunities.
This work addresses these challenges by focusing on optimizing HLS for FPGA design, with a particular emphasis on reducing latency while adhering to resource constraints. The goal is to transform affine high-level code (e.g., Listing 1) into an SLR-aware HLS-C++ design that achieves high QoR and facilitates automated bitstream generation. Achieving this requires addressing multiple challenges simultaneously, including task-to-SLR mapping, balancing computation and communication, and efficiently regenerating code to adapt to resource conflicts or spatial congestion. Existing methods fall short of providing an integrated solution capable of achieving these objectives, highlighting the need for a unified approach.
To address these limitations, we propose Prometheus, an integrated framework that combines the benefits of shared buffer and dataflow paradigms to optimize HLS-based FPGA design. The key contributions of this work include:
@@
  • A unified design space supporting tiled dataflow architectures, task fusion, tiling strategies, loop permutation, overlapping computation and communication, and concurrent task execution.
  • The application of a Non-Linear Problem (NLP) formulation to discover optimal design parameters within strict resource constraints, enabling efficient bitstream generation.
  • An end-to-end automated framework that generates OpenCL host code for correctness verification, performs Vitis HLS simulations, and produces the final bitstream.
  • Rigorous evaluation through RTL simulation and on-board evaluation, demonstrating superior QoR compared to state-of-the-art tools like AutoDSE [66], Sisyphus [59], and ScaleHLS [78].
@@ II-A HLS Optimization
HLS optimizations rely heavily on directives such as unroll, pipeline, and array_partitioning, which play a pivotal role in enhancing performance. The unroll pragma duplicates the loop body in the RTL design to enable parallel execution of loop iterations, while pipeline overlaps loop operations, reducing latency based on the initiation interval (II). Similarly, array_partitioning divides arrays into on-chip memory banks, allowing simultaneous access and facilitating more effective unrolling.
Advanced optimizations also leverage shared buffers to reduce off-chip memory transfers, ping-pong buffering to overlap data transfer and computation, and coarse-grained parallelism through duplicated processing elements [19]. Modern AMD/Xilinx FPGAs support data transfers with bit widths up to 512 bits, determined by problem dimensions. For instance, transferring 216 floats using a 256-bit width (8 floats per transfer) requires 27 transfers.
HLS tools like Vitis HLS support the dataflow pragma, which organizes computations into independent actors communicating through FIFO queues [40]. This structure enables concurrent execution by allowing each actor to process inputs and generate outputs as soon as data becomes available. Performance evaluations often combine estimation (e.g., Vitis HLS reports), RTL simulation, and on-board evaluation. While estimations and simulations avoid place-and-route overheads, on-board evaluations impose stricter constraints for generating deployable bitstreams.
II-B Limitations of Existing Methods
The limitations of current HLS optimization frameworks are illustrated through the 3mm kernel, which performs two independent matrix multiplications followed by a dependent multiplication. As shown in Listing 1, each matrix multiplication involves nested loops that can benefit from parallelism, unrolling, and communication-computation overlap.
@@ 15 G[i][j] += E[i][k] * F[k][j];} // S5
Shared Buffer  Frameworks such as AutoDSE [66], Sisyphus [59], ScaleHLS [78], and others [65, 63, 58, 57] optimize 3mm primarily through sequential execution of matrix multiplications. This approach enables reuse of DSP resources but fails to leverage opportunities for concurrent execution, such as running the first two multiplications in parallel. Current frameworks transfer arrays at different loop levels to manage on-chip memory but lack automatic mechanisms to overlap computation with communication effectively.
Although shared buffering could support concurrent execution of independent tasks, its application remains limited. Extending concurrency to all three multiplications would require task fusion, which may constrain compiler optimizations.
Dataflow  The dataflow paradigm, as implemented in frameworks like StreamHLS [9], divides computations into independent tasks, with communication managed through FIFOs. For the 3mm kernel in Listing 1, dataflow enables concurrent execution of the first two matrix multiplications, while the third begins as soon as its inputs are ready. This approach overlaps computation and communication, effectively managing task dependencies. As 3mm is computation-bound, increasing parallelism is essential. However, parallelism in pure dataflow methods is constrained by the FIFO bit width of the receiving task and the number of FIFOs. Increasing FIFO bit widths or adding more FIFOs to handle higher concurrency often results in routing congestion, making it challenging to generate a valid bitstream. Additionally, the lack of shared memory prevents data reuse across tasks, leading to repeated off-chip transfers or redundant buffers.
Mixed Methods  Mixed approaches, such as those used by Allo [15] and HIDA [77], combine shared buffering and dataflow paradigms. Allo extends HeteroCL [37] within the MLIR framework [39], supporting code transformation and dataflow scheduling. However, these methods lack automatic data-tile selection and pragma insertion, limiting their ability to optimize QoR effectively.
General Limitations  Most existing HLS frameworks fail to address key challenges such as automatic ping-pong buffering, tile size adjustments, and padding to maximize bit widths for communication efficiency. Unroll factors are often restricted to evenly divide problem dimensions, which can lead to inefficiencies with partial tiles. Expanding these factors and incorporating dynamic adjustments could significantly enhance parallelism and performance.
@@ II-C Overview of Prometheus
We present Prometheus, a unified optimization framework designed to efficiently explore design spaces using a Non-Linear Problem (NLP) cost model. Prometheus simplifies the development of FPGA designs by offering an engineer-friendly interface with adjustable parameters such as tile size, array partitioning, and unroll factors. It is SLR-aware, ensuring effective task distribution across Super Logic Regions (SLRs), which simplifies bitstream generation while optimizing the balance between performance and resource utilization.
The pseudo-code in Listing 2 demonstrates how Prometheus processes the 3mm kernel. Load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFO.
Each taski in the pseudo-code corresponds to the fully unrolled computation of an intra-tile for statement Si in Listing 1. An example of such a task is illustrated in Listing 3.
Arrays $E$, $F$, and $G$ are initialized to zero within their respective tasks (e.g., S0, S2, and S4) and are not preloaded, reducing unnecessary memory overhead.
Prometheus enhances computation efficiency by fusing statements that produce the same outputs (e.g., in Listing 1), and it automatically formulates an NLP problem to determine the theoretically optimal parameters, such as loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding. The framework dynamically adjusts bit widths for arrays like $F$, $D$, and $G$ to achieve a better balance between parallelism and resource usage.
@@
7 task0(E); // S0 + intra-tile loops
@@
10 task1(E, A, B); // S1 + intra-tile loops
@@
18 task2(F); // S2 + intra-tile loops
@@
21 task3(F, C, D); // S3 + intra-tile loops
@@
31 task4(G); // S4 + intra-tile loops
@@
34 task5(G, E, F); // S5 + intra-tile loops
@@
Efficient computation-communication overlap is achieved through ping-pong buffering, while concurrent task execution and FIFO-triggered dependent tasks maximize overall performance. Fused Tasks 0 and 1 execute concurrently, while Fused Task 2 begins as soon as the data tiles for $F$ and $E$ become available.
For the 3mm kernel optimized with RTL simulation, Prometheus achieved a throughput of 368.36 GF/s, significantly surpassing state-of-the-art frameworks such as Sisyphus (178.97 GF/s), Allo (60.40 GF/s), ScaleHLS (43.04 GF/s), and AutoDSE (1.74 GF/s). These results demonstrate the effectiveness of Prometheus in overcoming the limitations of existing methods and achieving superior QoR.
III Code Transformation
@@
Dependency Graph Creation  Our process starts with affine C/C++ code as input, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop. ISCC [68] verifies the legality of these transformations, ensuring dependencies are preserved. After achieving full distribution, we construct a dependency graph using PoCC [53]. PoCC provides the necessary information about the schedule and dependencies, enabling us to build the graph. In this graph, the nodes represent tasks, while the edges capture data communication arising from inter-task dependencies. Tasks with identical outputs are then merged (when legal), creating fused tasks with output-stationary properties. This ensures that each tile’s output is handled (loaded, computed, and either stored or transmitted) only once. If a dependency prevents distribution, the framework will still function but with a more limited optimization space, making it more effective when distribution is possible.
@@
Data-tiling and Padding  In the fused dependency graph, edges represent data communication between tasks as well as between tasks and off-chip memory, involving the transfer of data tiles with specified sizes. Data tiling divides loop iterations into smaller tiles, splitting array accesses into subsets. Each loop $l$ iterating over an array $a$ is divided into an outer loop ($TC^{l}_{inter}$) and an inner loop ($TC^{l}_{intra}$), with feasible permutations applied. For each array and dimension iterated by the loop $l$, the tile factor is a common choice that influences all arrays iterated by this loop.
@@
To optimize memory transfers, padding is applied to arrays in two ways. Simple padding increases the bit width ($BW_{a}$) for efficient transfers while maintaining the original loop trip count. Composite padding adjusts both $BW_{a}$ and the loop trip count ($TC^{l}$) to support unroll factors that do not evenly divide the original trip count, allowing irregular tile sizes and expanding the design space. Tile sizes are consistent within a fused task but vary between tasks. In Listing 2, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
@@
Fine-grained Parallelism  Data-tile selection involves splitting each loop and permuting them to create two levels of the original loops. Using ISCC [68], we verify the legality of these permutations, resulting in two loop levels: inter-tile (outer) and intra-tile (inner). If permutation is not feasible, the inter-tile loop retains its legal position. For tasks belonging to the same fused task, we merge their inter-tile loops, which are non-reduction. For instance, in Listing 2, the inter-tile loops on lines 3 and 5 iterate over task0 and task1. The intra-tile loops are fully unrolled, ensuring that all data accesses within the intra-tile remain on-chip. Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access. Reduction loops across inter-tiles are pipelined with an initiation interval ($II=n>1$), where n matches the reduction latency. For instance, in Listing 2, additions take 3 cycles, resulting in $II=3$.
@@
Loop Order  Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile. We place the inter-tile reduction loops directly above the task and are pipelined. If multiple reduction loops exist, we rank them by the size of their trip counts, placing the loop with the highest trip count innermost to ensure an efficient pipeline. This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops. This order will subsequently be determined by the NLP (cf. Section IV). As shown in Listing 2, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 1.
@@
Automatic Overlapping of Communication and Computation  To overlap communication and computation, we use on-chip buffers. The buffer size and data transfer location are determined by various options explored via NLP (cf. Section IV). Since data must stay on-chip for intra-task computation, transfers occur either below an inter-tile loop or before any loops start. The transfer position determines the data tile size, covering all data accessed below it. To improve data reuse, buffer size can match or exceed the transferred data tile. Similar to transfer location, buffer size is determined by its position relative to inter-tile loops or before any loops. Two boolean variables, $d_{a,l}$ and $t_{a,l}$, define and transfer array $a$ under loop $l$ when set to true. If defined or transferred before loops, $l=0$. Double-buffering is used for read-only or write-only arrays, and triple-buffering for arrays that are both read and written. Following [19], we perform an initial load, then overlap loading the next tile with computing the current one.
@@
Memory Transfer  Communication latency is reduced by transferring more data per cycle using increased bit width with padding. Read-only arrays are duplicated in off-chip memory for tasks with multiple reads, eliminating feed-through logic. For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory transfer.

### v3

@@ Abstract
Customized accelerators have transformed modern computing by enhancing energy efficiency and performance through specialization. Field Programmable Gate Arrays play a pivotal role in this domain due to their flexibility and high-performance potential. High-Level Synthesis and source-to-source compilers simplify hardware design by converting high-level code into hardware descriptions enriched with directives. However, achieving high Quality of Results in FPGA designs remains challenging, requiring complex transformations, strategic directive use, and efficient data management. While existing approaches like Design Space Exploration (DSE) and source-to-source compilers have made strides in improving performance, they often address isolated aspects of the design process.
This paper introduces Prometheus, a holistic framework that integrates task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution into a unified design space. Leveraging Non-Linear Problem methodologies, Prometheus explores this space to find solutions under resource constraints, enabling bitstream generation.
@@ I Introduction
The emergence of customized accelerators has revolutionized modern computing by delivering significant energy efficiency and performance gains through specialization. Field-Programmable Gate Arrays (FPGAs) have gained prominence for their flexibility and high-performance potential. High-level synthesis (HLS) and source-to-source compilers [19, 81, 76, 32, 62, 50, 77, 31, 79, 38, 80, 43] have made hardware design more accessible by converting high-level code (e.g., C++, Python) with hardware-specific directives into hardware descriptions. However, their effectiveness heavily relies on the quality of the input code, the applied hardware directives, and the use of advanced code transformations and directive strategies to optimize data communication and achieve high Quality of Result (QoR).
Challenges in Existing Approaches  Several works have contributed to enhancing performance through Design Space Exploration (DSE) [58, 59, 67, 66, 64, 24, 82, 66] and source-to-source compilers [38, 79, 78, 60, 71]. Despite these advancements, significant challenges remain, notably in supporting Code Transformation, which involves efficiently optimizing code to fully exploit FPGA potential; Overlapping Communication and Computation, aimed at maximizing concurrency to enhance performance; Concurrent Task Execution, which ensures simultaneous execution of tasks without resource conflicts; and Exploiting Intra-Task Parallelism, leveraging parallel execution within tasks to further boost performance. Previous works address specific optimization challenges but do not tackle all aspects simultaneously, underscoring the need for comprehensive, integrated solutions. Moreover, generating synthesizable code that can be converted into a bitstream without human intervention is a persistent challenge in FPGA design. This complexity is exacerbated by the intricate architecture of FPGAs, which often incorporates multiple Super Logic Regions (SLRs). In addition to adhering to the overall resource constraints of the entire FPGA board, it is crucial to meet specific constraints of each individual SLR while the following factors need to be considered to generate a high-quality HLS solution.
Shared Buffer  A classical approach to improving performance involves using shared buffers, which cache data on-chip to facilitate reuse by various computational tasks. This method increases parallelism and is particularly effective for computation-bound kernels. Studies have shown the effectiveness of shared buffers [66, 64, 79, 77, 38, 31, 60, 59, 58]. However, on-chip memory limitations often prevent caching all necessary data, and overlapping computation and communication remains constrained. While some compilers like Merlin [77] and PASTA [35] support double buffering for single arrays, Merlin’s application is limited and context-specific, and PASTA lacks parallelism.
Dataflow  On FPGAs, dataflow enables seamless data movement between off-chip and on-chip memory and computational tasks. This method supports the overlap of communication and computation, enabling concurrent execution of tasks and allowing dependent tasks to proceed once data is available via FIFO (First-In-First-Out). Dataflow is highly effective for memory-bound kernels but limits parallelism due to the absence of shared memory among tasks. Each task is compiled independently, which simplifies the compilation process and can enhance operating frequency [26, 27, 28].
Integrating shared buffers with dataflow methods combines their strengths, boosting performance and reducing resource use. A specialized design space for automated bitstream synthesis streamlines the process by reducing manual rewriting of non-synthesizable code, requiring efficient global and per-task resource management.
Prometheus Approach  To address these limitations and advance high-level source-to-source compilation, we propose Prometheus. This framework integrates the benefits of shared buffers and dataflow, aiming to improve performance while managing resources efficiently to ensure bitstream generation. The main contributions are:
@@
  • The introduction of a unified design space facilitates the development of a tiled dataflow architecture. This space incorporates task fusion, tiling strategies, loop permutation, overlapping of computation and communication automatically, and concurrent task execution.
  • Employing Non-Linear Problem methodology to navigate and discover theoretically optimal solutions within stringent resource constraints for generating bitstreams.
  • Automating our framework that is end-to-end comprehensive. It automatically generates a host file in OpenCL to verify correctness, runs Vitis HLS for report generation and RTL simulations, and generates the bitstream.
  • Conducting evaluations utilizing RTL simulation and on-board FPGA runs, delivering superior Quality of Results (QoR) compared to AutoDSE [67], Sisyphus [60], ScaleHLS [79], and Allo [14]. Importantly, it achieves these results without requiring DSE for HLS design generation.
@@ II-A HLS optimization
In High-Level Synthesis (HLS), directives such as unroll, pipeline, and array_partitioning pragmas play a crucial role. The unroll pragma enables parallel execution by duplicating statements, while pipeline stages loops for overlapping operations based on the initiation interval (II). array_partitioning divides arrays into on-chip memory banks, facilitating simultaneous data access and enhancing unrolling. Additional optimizations include shared buffers for reducing off-chip transfers, coarse-grained parallelism through duplicated processing elements, and ping-pong buffering to overlap data transfer and computation [18]. On AMD/Xilinx FPGAs, data transfers use bit widths up to 512 bits, with the maximum bit width determined by problem dimensions. For instance, transferring 216 floats with a 256-bit width (8 floats per transfer) requires 27 transfers. The dataflow pragma organizes computation into actors that communicate via FIFO queues, where each actor processes inputs and produces outputs as they become available [41].
Performance evaluation methods include estimation (e.g., Vitis HLS reports), RTL simulation, and on-board evaluation. Estimation and simulation avoid place-and-route but utilize board resources, while on-board evaluation imposes stricter constraints for bitstream generation.
II-B Limitation of existing methods
To demonstrate the limitations of the existing framework and methodology we use 3mm. As illustrated in Listing 1, two independent matrix multiplications compute matrices that are then utilized in the final matrix multiplication.
Shared Buffer  AutoDSE [67], Sisyphus [60], ScaleHLS [79], and frameworks like [66, 64, 59, 58] optimize the 3mm kernel by sequentially executing each matrix multiplication. AutoDSE uses a bottleneck-based DSE approach, Sisyphus employs loop splitting and NLP-based cost modeling, and ScaleHLS automates multi-level optimizations in Vitis HLS.
Sequential execution enables DSP reuse but lacks support for concurrent task execution, e.g., the first two matrices in 3mm are independent and can run concurrently. Although arrays can be transferred at different loop levels to manage on-chip memory, automatic computation-communication overlap is not achieved. Current frameworks enhance performance by transferring large arrays on-chip and extensively unrolling loops to maximize parallelism.
Concurrent execution of the first two matrix multiplications could be achieved using a shared buffering method, but this remains unexplored. Extending concurrency to all three operations would require fusion techniques, potentially limiting the compiler’s optimization capabilities.
@@ 15 G[i][j] += E[i][k] * F[k][j];} // S5
Dataflow  The dataflow method creates six tasks, each corresponding to a statement, with communication managed through FIFOs. For the 3mm kernel, efficient data reuse is critical and can be achieved via reuse buffers or repeated off-chip data transfers. However, reuse buffers are limited to individual tasks and cannot be shared across tasks.
As 3mm is computation-bound, increasing parallelism is vital. However, parallelism is constrained by the receiving task’s FIFO bit width, and attempts to scale this approach often lead to routing congestion and impractical bitstream generation. Despite these challenges, dataflow allows concurrent execution of the first two matrix multiplications and starts the third as soon as data becomes available.
Mixed Method  Methods like Allo [14] and HIDA [78] integrate shared buffers and dataflow techniques to enhance QoR. Allo extends HeteroCL [38] within the MLIR framework [40], supporting code transformation and dataflow but lacks automatic data-tile selection and pragma insertion.
Limitations in General  Existing HLS frameworks often lack key optimizations, such as automatic ping-pong buffering for overlapping computation and communication, tile size adjustments, and padding to maximize bit width for reducing communication latency. These gaps can lead to suboptimal performance, particularly in data-intensive scenarios. Most frameworks [60, 59] restrict unroll factors to evenly divide problem sizes, while others [67, 66] leveraging Merlin [77] permit uneven unroll factors. However, optimizations typically target complete tiles, resulting in inefficiencies with partial tiles and non-innermost unrolled loops. Expanding unroll factor options could better balance parallelism and performance.
For the 3mm kernel optimized with RTL simulation, our framework (Prometheus) achieved a throughput of 368.36 GF/s, significantly surpassing Sisyphus (178.97), Allo (60.40), ScaleHLS (43.04), and AutoDSE (1.74). These frameworks employ varying strategies—AutoDSE, Sisyphus, and ScaleHLS use shared buffer methods, while Allo applies a mixed method. However, due to the limitations noted, further improvements are needed to achieve a superior QoR.
@@ II-C Overview of Prometheus
We propose Prometheus, a unified optimization framework leveraging a Non-Linear Problem (NLP) cost model to explore design spaces efficiently. Prometheus is engineer-friendly, offering adjustable parameters such as tile size, array partitioning, and unroll factors, while being SLR-aware for task distribution across SLRs, simplifying bitstream generation and optimizing performance-resource balance.
The pseudo-code in Listing 2 illustrates how Prometheus processes the 3mm kernel. The load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFOs. Each task_i represents the fully unrolled computation of the intra-tile for statement Si in Listing 1. We can see an example on Listing 3. Arrays E, F, and G are initialized to 0 with statements S0, S2, and S4, respectively, and are not loaded. Prometheus optimizes computation by fusing statements producing the same outputs (e.g., Listing 1), automatically generating an NLP problem to determine loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding. It strategically adjusts bit widths (e.g., for arrays F, D, and G) to balance parallelism and resource usage.
@@
7 task_0(E); // S0 + intra-tile loops
@@
10 task_1(E, A, B); // S1 + intra-tile loops
@@
18 task_2(F); // S2 + intra-tile loops
@@
21 task_3(F, C, D); // S3 + intra-tile loops
@@
31 task_4(G); // S4 + intra-tile loops
@@
34 task_5(G, E, F); // S5 + intra-tile loops
@@
Efficient overlap of computation and communication is achieved via ping-pong buffering, concurrent task execution, and FIFO-triggered dependent tasks. Fused tasks 0 and 1 execute concurrently, and Task 2 starts once arrays F and E are ready. These optimizations enable Prometheus to surpass frameworks like Sisyphus [60] in throughput and QoR, particularly for computation-bound kernels. Its SLR-awareness and dataflow integration make it practical for generating deployable, resource-efficient bitstreams.
III Problem Formulation
@@
Dependency Graph Creation  Our process begins with the input C/C++ code, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop. ISCC [69] verifies the legality of these transformations, ensuring dependencies are preserved. After achieving full distribution, where each loop body corresponds to a single task, we construct a dependency graph capturing inter-task dependencies. Tasks with identical outputs are then merged (when legal), creating fused tasks with output-stationary properties. If a dependency prevents distribution, the framework will still function but with a more limited optimization space, making it more effective when distribution is possible.
@@
Data-tiling and Padding  From the fused dependency graph, edges represent data communication between tasks and off-chip memory, transferring data-tiles of specified sizes. Data tiling divides loop iterations into smaller tiles, splitting array accesses into subsets. Each loop $l$ iterating over an array $a$ is divided into an outer loop ($TC^{l}_{inter}$) and an inner loop ($TC^{l}_{intra}$), with feasible permutations applied. If an array is not fully transferred in a dimension $n$, its tile size is constrained by other arrays sharing the same loop. To optimize memory transfers, padding is applied to arrays in two ways. Simple padding increases the bit width ($BW_{a}$) for efficient transfers while maintaining the original loop trip count. Composite padding adjusts both $BW_{a}$ and the loop trip count ($TC^{l}$) to support unroll factors that do not evenly divide the original trip count, allowing irregular tile sizes and expanding the design space. Tile sizes are consistent within a fused task but vary between tasks. For example, in Listing 2, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
@@
Fine-grained Parallelism  Data-tile selection involves splitting each loop and permuting the outermost loop. Using ISCC [69], we ensure the legality of permutations, yielding two loop levels: inter-tile (outer) and intra-tile (inner). If permutation is infeasible, the inter-tile loop remains in its legal position. In Listing 2, inter-tile loops (lines 3 and 5) iterate intra-tile loops in task_0, which are fully unrolled to maximize parallelism. Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access. Reduction loops across inter-tiles are pipelined with an initiation interval ($II=n>1$), where n matches the reduction latency. For instance, in Listing 2, additions take 3 cycles, resulting in $II=3$. Fused tasks are output-stationary, ensuring each tile’s output is processed (loaded, computed, and stored or sent) only once. Dependent tasks execute sequentially with high throughput, leveraging unrolled tasks and on-chip data. Processed data-tiles are then transferred to the next task or off-chip memory, optimizing performance.
@@
Loop Order  Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile. We place the inter-tile reduction loops directly above the task and are pipelined. If multiple reduction loops exist, we rank them by the size of their trip counts, placing the loop with the highest trip count innermost to ensure an efficient pipeline. This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops. This order will subsequently be determined by the NLP, as described in Section IV. As shown in Listing 2, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 1.
@@
Automatic Overlapping of Communication and Computation  To overlap communication and computation, we use on-chip buffers. The buffer size and data transfer location are determined by various options explored via NLP (cf. Section IV). Since data must stay on-chip for intra-task computation, transfers occur either below an inter-tile loop or before any loops start. The transfer position determines the data tile size, covering all data accessed below it. To improve data reuse, buffer size can match or exceed the transferred data tile. Similar to transfer location, buffer size is determined by its position relative to inter-tile loops or before any loops. Two boolean variables, $d_{a,l}$ and $t_{a,l}$, define and transfer array $a$ under loop $l$ when set to true. If defined or transferred before loops, $l=0$. Double-buffering is used for read-only arrays, and triple-buffering for arrays that are both read and written. Following [18], we perform an initial load, then overlap loading the next tile with computing the current one.
@@
Memory Transfer  Communication latency is reduced by transferring more data per cycle using increased bit width with padding. Read-only arrays are duplicated in off-chip memory for tasks with multiple reads, eliminating feed-through logic. For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory communication.